OpenStudy (anonymous):

Use the Pythagorean Theorem to find the length of the missing side of the right triangle with both legs equal to √2. a right triangle with legs as square root of 2 and square root of 2. A. 2 B. 2 √2 C. √5 D. None of the above

OpenStudy (anonymous):

@Fugazi

OpenStudy (anonymous):

honestly do not understand this ))): sorry!!!!

OpenStudy (anonymous):

@Fugazi its totally ok. i don't either but thanks for the help

OpenStudy (anonymous):

Hey,apply the pythagoras theorem!

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (anonymous):

Yup,thts ri8!

OpenStudy (anonymous):

what do i do now?

OpenStudy (anonymous):

Here its given b=c=sqrt(2)

OpenStudy (anonymous):

So we have to find a

OpenStudy (anonymous):

sorry Its a=b=sqrt{2}

OpenStudy (anonymous):

its c we have to find :)

OpenStudy (anonymous):

ooooh.√(2^2) = 2 so 2+2= 4 √(4)= c^2 2=c

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

You are good to go!Nice work!:)

OpenStudy (anonymous):

literally thank you soon much!